An element with mass 640 grams decays by 7.3% per minute. How much of the element is remaining after 8 minutes, to the nearest 10th of a gram?



1
Expert's answer
2022-03-25T14:50:26-0400

Solution

The exponential decay equation is given by :-

"y=A(1-r)^{t}"

 Where

A= Initial value ,

r = rate of decay in decimal and t is time.

As per given , we have

A= 640 gm

r= 7.3%=0.073

Then , our decay function will become :-

"y=640(1-0.073)^{t}"

"y=640(0.927)^{8}"

"y=640\times0.54530298264"

y=348.99

y=349.0

At t= 8 minutes, we have


Hence, after 8 minutes , the remaining amount of element = 349.0 gms.